The sum of three numbers is 20. If we multiply the first number by 2 and add the second

number and subtract the third number, then we get 23. If we multiply the first number by 3

and add second and third number to it, then we get 46. Let x be the first number, y be the

second number and z


be the third number.


(a) Obtain a system of linear equations to represent the given information.

(b) Write down the system in (a) as a matrix equation.

(c) Use inverse matrix to solve for x , y and z .


1
Expert's answer
2021-09-07T17:32:59-0400

(a) Obtain a system of linear equations to represent the given information.

Construct a system containing three linear equations.


"x+y+z=20"

"2x+y-z=23"

"3x+y+z=46"

 (b) Write down the system in (a) as a matrix equation.


"AX=B"

"A=\begin{pmatrix}
 1 & 1 & 1\\
 2 & 1 & -1\\
 3 & 1 & 1\\
\end{pmatrix}, X=\begin{pmatrix}
 x\\
 y \\
 z
\end{pmatrix}, B=\begin{pmatrix}
 20\\
 23\\
46
\end{pmatrix}"

"\begin{pmatrix}
 1 & 1 & 1\\
 2 & 1 & -1\\
 3 & 1 & 1\\
\end{pmatrix}\begin{pmatrix}
 x\\
 y \\
 z
\end{pmatrix}=\begin{pmatrix}
 20\\
 23\\
46
\end{pmatrix}"

(c) Use inverse matrix to solve for "x,y" and "z"


"AX=B"

"A^{-1}AX=A^{-1}B"

"X=A^{-1}B"

"\\det\u2061A=\\begin{vmatrix}\n 1 & 1 & 1\\\\\n 2 & 1 & -1\\\\\n 3 & 1 & 1\\\\\n\\end{vmatrix}=1\\begin{vmatrix}\n 1 & -1 \\\\\n 1 & 1\n\\end{vmatrix}-1\\begin{vmatrix}\n 2 & -1 \\\\\n 3 & 1\n\\end{vmatrix}+1\\begin{vmatrix}\n 2 & 1 \\\\\n 3 & 1\n\\end{vmatrix}"

"=1+1-(2+3)+2-3=-4\not=0=>A^{-1}\ exists"

Find the cofactor matrix:


"C_{11}=(-1)^{(1+1)}\begin{vmatrix}
 1 & -1 \\
 1 & 1
\end{vmatrix}=2"

"C_{12}=(-1)^{(1+2)}\begin{vmatrix}
 2 & -1 \\
 3 & 1
\end{vmatrix}=-5"

"C_{13}=(-1)^{(1+3)}\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}=-1"

"C_{21}=(-1)^{(2+1)}\begin{vmatrix}
 1 & 1 \\
 1 & 1
\end{vmatrix}=0"


"C_{22}=(-1)^{(2+2)}\begin{vmatrix}
 1 & 1 \\
 3 & 1
\end{vmatrix}=-2"


"C_{23}=(-1)^{(2+3)}\begin{vmatrix}
 1 & 1 \\
 3 & 1
\end{vmatrix}=2"

"C_{31}=(-1)^{(3+1)}\begin{vmatrix}
 1 & 1 \\
 1 & -1
\end{vmatrix}=-2"

 

"C_{32}=(-1)^{(3+2)}\begin{vmatrix}
 1 & 1 \\
 2 & -1
\end{vmatrix}=3"

"C_{33}=(-1)^{(3+3)}\begin{vmatrix}
 1 & 1 \\
 2 & 1
\end{vmatrix}=-1"

The cofactor matrix is


"\begin{pmatrix}
 2 & -5 & -1\\
 0 & -2 & 2\\
 -2 & 3 & -1\\
\end{pmatrix}"

The adjugate matrix is


"\begin{pmatrix}
 2 & 0 & -2\\
 -5 & -2 & 3\\
 -1 & 2 & -1\\
\end{pmatrix}"


"A^{-1}=\dfrac{1}{-4}\begin{pmatrix}
 2 & 0 & -2\\
 -5 & -2 & 3\\
 -1 & 2 & -1\\
\end{pmatrix}"

"=\begin{pmatrix}
 -1/2 & 0 &1/2\\
 5/4 & 1/2 & -3/4\\
 1/4 & -1/2 & 1/4\\
\end{pmatrix}"

"\begin{pmatrix}
 x\\
 y \\
 z
\end{pmatrix}=\begin{pmatrix}
 -1/2 & 0 &1/2\\
 5/4 & 1/2 & -3/4\\
 1/4 & -1/2 & 1/4\\
\end{pmatrix}\begin{pmatrix}
 20\\
 23\\
46
\end{pmatrix}"


"=\begin{pmatrix}
 -10+0+23 \\
 25+23/2-69/2 \\
 5-23/2+23/2\\
\end{pmatrix}=\begin{pmatrix}
 13\\
 2\\
 5\\
\end{pmatrix}"

Then "x=13, y=2, z=5."

 

"(13, 2, 5)."